Assume 20 degrees. What is the wavelength of light passing through double slit that creates a pattern on a screen 2 meters away. Slits are seperated by 2.50*10^-3. You detect third dark fringe at 2.50*10^-2 above central bright fringe.
Answer #1

Solution y = (261) (2h-1) do 2 d for darle dringes Da 2m d - 2.5X10-3m. G = 2.5x10-2m so) h=3 2 2-5X10? = (3-k) x + x2 2.5810